If the total bill at a resturant, including a 15% tip, is $65.32, what was the cost of the meal alone?
Answer by tutorcecilia(2152)

(Cost of Meal)+(Cost of Meal)(tax)=$65.32 [plug-in the values]
C+.15C=65.32 [Change 15% to .15; solve for C]
1.15C=65.32
1.15C/1.15=65.32/1.15
C=$56.80
.
Check by plugging all values back into the orginal equation and solve:
C+ .15C=65.32
(56.80)+ .15(56.80)=65.32
65.32=65.32 [checks out]